OpenStudy (anonymous):

Help!! Simplify: [((√(x^2+1)-x) x (2x/(2√(x^2+1))]/(x^2+1)

OpenStudy (anonymous):

\[(\sqrt{x^2+1}-x \times (2x/2\sqrt{x^2+1}))/\sqrt{x^2+1}\]

OpenStudy (mathstudent55):

Is this it? \(\dfrac{(\sqrt{x^2 + 1} - x)\frac{2x}{2 \sqrt{x^2 + 1} }}{x^2 + 1} \)

OpenStudy (anonymous):

Yeah

OpenStudy (mathstudent55):

\(=\dfrac{(\sqrt{x^2 + 1} - x)\frac{\cancel{2}x}{\cancel{2} \sqrt{x^2 + 1} + 1}}{x^2 + 1} \) \(= \dfrac{(\sqrt{x^2 + 1} - x)\frac{x}{ \sqrt{x^2 + 1} + 1}}{x^2 + 1} \) \(= \dfrac{x(\sqrt{x^2 + 1} - x) }{(x^2 + 1)( \sqrt{x^2 + 1} + 1 )} \)

OpenStudy (anonymous):

where did the 1+1 come from??

OpenStudy (mathstudent55):

You're right. I made a mistake when I copied one step to the next.

OpenStudy (mathstudent55):

\(=\dfrac{(\sqrt{x^2 + 1} - x)\frac{\cancel{2}x}{\cancel{2} \sqrt{x^2 + 1}}}{x^2 + 1}\) \(=\dfrac{(\sqrt{x^2 + 1} - x)\frac{x}{\sqrt{x^2 + 1}}}{x^2 + 1}\) \(=\dfrac{x(\sqrt{x^2 + 1} - x)} {(x^2 + 1) \sqrt{x^2 + 1} }\) Now we can rationalize the denominator: \(=\dfrac{x(\sqrt{x^2 + 1} - x)} {(x^2 + 1) \sqrt{x^2 + 1} } \times \dfrac{ \sqrt{x^2 + 1} }{ \sqrt{x^2 + 1} }\) \(=\dfrac{x \sqrt{x^2 + 1} (\sqrt{x^2 + 1} - x)} {(x^2 + 1) ^2 }\)